# Multi-student Diffusion Distillation for Better One-step Generators

### v1

@@ 5.4 Ablation Studies
Table 3: Ablation studies on different components of MSD. All experiments are done on ImageNet-64$\times$64, trained with only the DM stage for $20$k iterations, where $B$ is the batch size per student. See App. A.1.
@@
MSD is still better with the same effective batch size. To investigate if the performance boost from MSD comes from only a batch size increase over single student distillation, we make a comparison with the same effective batch size. As showcased in Tab. 3, MSD with $4$ students and a batch size of $32$ per student performs slightly better than the single-student counterpart with a batch size of $128$, indicating that MSD likely benefits from a capacity increase than a batch size increase. As a takeaway, with a fixed training resource measured in processed data points, users are better off distilling into multiple students with partitioned resources each than using all resources to distill into a single student. Although multiple students means multiple model weights to save, storage is often cheap, so in many applications, this cost is outweighed by our improved quality or latency.
Simple splitting works surprisingly well. We used consecutive splitting of classes in Sec.  5.2. In the pretrained EDM model, class labels are fed to the diffusion model via the label embedding layer. Therefore, we investigated another strategy where we performed a $K$-means clustering ($K=4$) on the label embeddings, resulting in $4$ clusters of similar sizes: $(230,283,280,207)$. However, MSD trained with these clustered partitions performs similarly to sequential partition, as shown in Tab. 3. For text-to-image generation, we performed $K$-means clustering on the pooled embeddings of prompts in the training data, resulting in clusters of vastly uneven sizes. Due to computational limitations, we opted for the simpler partition strategies outlined in Sec. 5.
Effect of scaling the number of students. In Tab. 3, we study the effect of increasing $K$, the number of distilled students. We kept the per-student batch size fixed so more students induce a larger effective batch size. We observe better FID scores for more students. We hypothesize that better training strategies, such as per-student tuning, will further improve the quality. Optimal strategies for scaling to ultra-large numbers of students is an interesting area for future work.
@@ 6.1 Limitations
MSD is the first work to explore diffusion distillation with multiple students, and it admits a few limitations that call for future work. 1) Further explorations could offer more insights into optimal design choices for a target quality and latency on various datasets, such as the number of students, input condition size for each student, and other hyperparameters. This is especially beneficial if the training budget is limited. 2) We apply simple partitioning for both class- and text-conditions and assign them disjointly to different students. Although our empirical study shows that simple alternatives do not offer obvious advantages, more sophisticated routing mechanisms may help. 3) We use simple channel reduction when designing smaller students to demonstrate feasibility. This results in a significantly smaller latency reduction than sample size reduction. Exploring other designs of smaller students will likely increase their quality and throughput. 4) We train different students separately, but we expect that carefully designed weight-sharing, loss-sharing, or other interaction schemes can further enhance training efficiency. 5) We hypothesize that MSD can be applied to other diffusion distillation methods and other modalities for similar benefits but leave this for future work.
@@ 6.2 Conclusion
This work presented Multi-Student Distillation, a simple yet efficient method to increase the effective model capacity for single-step diffusion distillation. We applied MSD to the distribution matching and adversarial distillation methods. We demonstrated their superior performance over single-student counterparts in class-conditional and text-to-image generations. Particularly, MSD with DMD2’s the two-stage training achieves state-of-the-art FID scores. Moreover, we successfully distilled smaller students from scratch, demonstrating MSD’s potential in further reducing the generation latency with multiple smaller student distillations. We envision building on MSD to enable generation in real-time, enabling many new use cases.
@@
Our work aims to improve the quality and speed of diffusion models, thus we may inherit ethics concerns from diffusion models and generative models in general. Potential risks include fabricating facts or profiles that could mislead public opinion, displaying biased information that may amplify social biases, and displacing creative jobs from artists and designers.
@@ Acknowledgements
Appendix A Additional ablation studies
A.1 Training curves for Sec. 5.4
@@ Appendix A Additional ablation studies
A.2 The effect of paired dataset size on DMD
@@
A.3 Single-step distillation methods comparison
@@
Method on ImageNet Terminal FID Iterated Images Note
@@
Table 5 justifies our choice of DMD/DMD2 as our first-stage training without adversarial loss. Competitor methods suffer from training instability (SiD, FID fluctuates more than 100%), larger training data size (EMD), or worse quality (CTM and other CM-based methods). DMD/DMD2, on the other hand, strike a good balance. We noticed DMD exhibits more stability for MSD on ImageNet, whereas DMD2 performs better for SD v1.5, which leads to our respective choices. As pointed out in App. B, we used a smaller-sized paired dataset (10 0001000010\,00010 000 images) than the original DMD paper (25 0002500025\,00025 000 images) for ImageNet, which significantly accelerated convergence without impairing the final performance. Moreover, as pointed out in Sec. 4.2, the same paired dataset can be used for all students, eliminating potential additional computation.
A.4 More results on distilling into smaller students
@@
Figure 8: The effect of student size, number of classes covered, and training stage on the performance of a single smaller student. XS, XXS, and B refer to model sizes, with precise definitions provided in Tab. 6. Also, see special evaluation details in App. C.
In Sec. 5.2, we trained MSD4-ADM on smaller students to demonstrate the tradeoff between generation quality and speed. Here, we make a more comprehensive ablation study on the interplay between student size, number of classes covered, and training stage, with results displayed in Fig. 8. We observe that generation quality increases with student size and decreases with more classes covered. MSD offers great flexibility for users to make these choices based on computational resources, generation quality, and inference speed requirements.
Appendix B Implementation Details
B.1 Toy Experiments
The real dataset is a mixture of Gaussians. The radius for the outer circle is 0.5, the radius for the 8 inner circles is 0.1, and the standard deviation for each Gaussian (smallest circle) is $0.005$. The teacher is trained with EDM noise schedule, where use $(\sigma_{\min},\sigma_{\max})=(0.002,80)$ and discretized the noise schedule into 1000 steps. We train the teacher for 100 000100000100\,000100 000 iterations with AdamW [38] optimizer, setting the learning rate at 1e-4, weight decay to 0.01, and beta parameters to (0.9, 0.999). For distillation, we first generated dataset of 1000 pairs, then used DMD [75] to train 1, 2, 4, and 8 students, respectively, all for 200 000200000200\,000200 000 iterations, with reduced learning rate at 1e-7. We only sample the first $750$ of the $1000$ steps for distillation.
@@ B.1 Toy Experiments
B.2 ImageNet
B.2.1 Same-sized students
@@
For the ADM stage, we attach a prediction head to the middle block of the fakescore model. The prediction head consists of a stack of $4\times 4$ convolutions with a stride of 2, group normalization, and SiLU activations. All feature maps are downsampled to $4\times 4$ resolution, followed by a single convolutional layer with a kernel size and stride of 4. The final output linear layer maps the given vector to a scalar predicted probability. We load the best generator checkpoint from DM stage, but re-initialize the fakescore and GAN classifier model from teacher weights, as we observed this leads to slightly better performance. We set the GAN generator loss weight to 3e-3 and the GAN discriminator loss weight to 1e-2, and reduce the learning rate to 5e-7. Each student is trained on 4 A100 GPUs, with a total batch size of 192, for 150 000150000150\,000150 000 iterations. This yields the MSD4-ADM checkpoint in Tab. 2.
B.2.2 Smaller students
@@
For the MSD4-ADM-S checkpoint in Tab. 2, we train the TSM stage using the model architecture S with the continuous EDM noise schedule with $(P_{\text{mean}},P_{\text{std}})=(-1.2,1.2)$ and the weighting $\lambda_{t}=(\sigma_{t}^{2}+0.5^{2})/(\sigma_{t}\cdot 0.5)^{2}$. We use a learning rate of 1e-4. Each student is trained on 4 A100 GPUs, with a total batch size of 576, for 400 000400000400\,000400 000 iterations. Then DM and ADM stages were trained using a total batch size of 160, following otherwise the same setup as Sec. B.2.1.
For the ablation study in A.4, we instead train a common TSM stage for all students for computational efficiency. We train this common stage using 16 A100 GPUs, with a total batch size of $3584$ and $4864$ for architecture XS and XXS, respectively. The DM and ADM stages are followed by specialized students with filtered data and 4 A100 GPUs each, with a total batch size of 224 and 256 for architecture XS and XXS, respectively, and using the same learning rate of 2e-6.
B.3 SD v1.5
B.3.1 Same-sized students, CFG=1.75
Our SD v1.5 experiments setup closely follows DMD2 [74] paper. We distill our one-step generators from the SD v1.5 [54] model, using a classifier-free guidance scale of $1.75$ for the teacher model to obtain the best FID score. We use the first 5M prompts from the COYO dataset [8], and the corresponding 5M images for the GAN discriminator. We apply the DDIM noise schedule with 1000 steps for sampling $t$. The weight $w_{t}$ in Eq. 2 is set to $\frac{\sigma_{t}^{2}}{\alpha_{t}}\frac{CS}{\|\bm{\mu}_{\text{teacher}}(\bm{x}_{t},t)-\bm{x}\|_{1}}$ where $S$ is the number of spatial locations and $C$ is the number of channels, and the weight $\lambda_{t}$ in Eq. 4 is set to $\alpha_{t}^{2}/\sigma_{t}^{2}$.
@@
For the ADM stage, we attach a prediction head that has the same architecture (though with a different input size) as Sec. B.2. We load both the best generator checkpoint and the corresponding fakescore checkpoint from DM stage. We set the GAN generator loss weight to 1e-3 and the GAN discriminator loss weight to 1e-2, and reduce the learning rate to 5e-7. Each student is trained on 32 A100 GPUs, with a total batch size of 1024, for $5000$ iterations. This yields the MSD4-ADM checkpoint in Tab. 2.
B.3.2 Same-sized students, CFG=8
@@ B.3.3 Smaller student, CFG=8
To create a subset of dog-related data, we first selected 1 210 000similar-toabsent1210000$\sim$\$1\,210\,000\$∼ 1 210 000 prompts in the COYO [8] dataset whose embeddings are closest to “a dog.” We then created an equal number of noise-image pairs from the SD v1.5 teacher using these prompts with CFG=8. We train the TSM stage using the model architecture S with the 1000-step DDIM noise schedule and the weighting $\lambda_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$. We use a learning rate of 1e-4. We then continue to the DM stage with the paired regression loss, using a learning rate of 1e-5, and finally continue to the ADM stage using generated paired images as ”real” images with a learning rate of 5e-7. We use 16 A100 GPUs. The TSM stage is trained with a total batch size of $1536$ for 240 000240000240\,000240 000 iterations. The DM stage is trained with a total batch size of $512$ for both the paired and the fake images for 20 0002000020\,00020 000 iterations, and the ADM stage is trained with the same batch size for $6000$ iterations. This yields the checkpoint used to generate Fig. 2(c). Longer training and better tuning are again likely to improve the generation quality further.
Appendix C Evaluation Details
For zero-shot COCO evaluation, we use the exact setup as GigaGAN [76] and DMD2 [74]. Specifically, we generate 30 0003000030\,00030 000 images using the prompts provided by DMD2 code. We downsample the generated images using PIL to $256\times 256$ Lanczos resizer. We then use the clean-fid [49] to compute the FID score between generated images and 40 5044050440\,50440 504 real images from the COCO 2014 validation dataset. For ImageNet, we generate 50 0005000050\,00050 000 images and calculate FID using EDM [25] evaluation code. When selecting the best checkpoints for partitioned students, the same procedure is applied only for prompts/classes within respective partitions. For the ablation study in Sec. A.4, 10 0001000010\,00010 000 images are generated for only the first $10$ classes for an apple-to-apple comparison.
Appendix D Additional qualitative results
D.1 Additional Imagenet-64$\times$64 results

### v2

@@ 5.4 Ablation Studies
Table 3: Ablation studies on different components of MSD. All experiments are done on ImageNet-64$\times$64, trained with only the DM stage for $20$k iterations, where $B$ is the batch size per student. See App. B.1.
@@
$4$ students, $B=128$, random splitting $2.45$
MSD is still better with the same effective batch size. To investigate if the performance boost from MSD comes from only a batch size increase over single student distillation, we make a comparison with the same effective batch size. As showcased in Tab. 3, MSD with $4$ students and a batch size of $32$ per student performs slightly better than the single-student counterpart with a batch size of $128$, indicating that MSD likely benefits from a capacity increase than a batch size increase. As a takeaway, with a fixed training resource measured in processed data points, users are better off distilling into multiple students with partitioned resources each than using all resources to distill into a single student. Our previous experiments also reflected this, where we used significantly fewer resources per student than the single-student counterparts (see details in App. D). Although multiple students means multiple model weights to save, storage is often cheap, so in many applications, this cost is outweighed by our improved quality or latency.
Simple splitting works surprisingly well. We used consecutive splitting of classes in Sec.  5.2. Although it shows an obvious advantage over random splitting, as shown in Tab. 3, it does not use the embedding information from the pretrained EDM model. Therefore, we investigated another strategy where we performed a $K$-means clustering ($K=4$) on the label embeddings, resulting in $4$ clusters of similar sizes: $(230,283,280,207)$. However, MSD trained with these clustered partitions performs similarly to sequential partition, as shown in Tab. 3. For text-to-image generation, we performed $K$-means clustering on the pooled embeddings of prompts in the training data, resulting in clusters of vastly uneven sizes. Due to computational limitations, we opted for the simpler partition strategies outlined in Sec. 5.
Effect of scaling the number of students. In Tab. 3, we study the effect of increasing $K$, the number of distilled students. We fixed the per-student batch size so that more students would induce a larger effective batch size. We observe better FID scores for more students. We hypothesize that better training strategies, such as per-student tuning, will further improve the quality. Optimal strategies for scaling to ultra-large numbers of students is an interesting area for future work.
@@ 6.1 Limitations
MSD is the first work to explore diffusion distillation with multiple students, and it admits a few limitations that call for future work. 1) Further explorations could offer more insights into optimal design choices for quality and latency on various datasets, such as the number of students, input condition size for each student, and other hyperparameters. This is especially beneficial if the training budget is limited. 2) We apply simple partitioning for both class- and text-conditions and assign them disjointly to different students. Although our empirical study shows that simple alternatives do not offer obvious advantages, more sophisticated routing mechanisms may help. 3) We use simple channel reduction when designing smaller students to demonstrate feasibility. This results in a significantly smaller latency reduction than the reduction in sample size. Exploring other designs of smaller students will likely increase their quality and throughput. 4) We train different students separately, but we expect that carefully designed weight-sharing, loss-sharing, or other interaction schemes can further enhance training efficiency. 5) We hypothesize that MSD can be applied to other state-of-the-art diffusion distillation methods and other modalities for similar benefits but leave this for future work.
@@ 6.2 Conclusion
This work presented Multi-Student Distillation, a simple yet efficient method to increase the effective model capacity for single-step diffusion distillation. We applied MSD to the distribution matching and adversarial distillation methods. We demonstrated their superior performance over single-student counterparts in class-conditional and text-to-image generations. Particularly, MSD with DMD2’s two-stage training achieves state-of-the-art FID scores. Moreover, we successfully distilled smaller students from scratch, demonstrating MSD’s potential in further reducing the generation latency with multiple smaller student distillations. We envision building on MSD to enable generation in real-time, enabling many new use cases.
@@
Our work aims to improve the quality and speed of diffusion models. Thus, we may inherit ethics concerns from diffusion and generative models. Potential risks include fabricating facts or profiles that could mislead public opinion, displaying biased information that may amplify social biases, and displacing creative jobs from artists and designers.
@@ Acknowledgements
Appendix A Additional experimental results
A.1 CLIP score for high guidance scale
Table 5: CLIP score comparison for high guidance scale on MS-COCO2014. LCM-LoRA is trained with a classifier-free guidance (CFG) scale of $7.5$, while other methods use a CFG of $8$.
Method Latency ($\downarrow$) CLIP score ($\uparrow$)
DPM++ (4 step) [40] $0.26$s $0.309$
UniPC (4 step) [80] $0.26$s $0.308$
LCM-LoRA (1 step) [42] $0.09$s $0.238$
LCM-LoRA (4 step) [42] $0.19$s $0.297$
DMD2 (our reimplementation) [75] $0.09$s $0.306$
MSD4-ADM (ours) $0.09$s $0.308$
DMD [76] $0.09$s $0.320$
SDv1.5 (teacher) [54] $2.59$s $0.322$
Tab. 5 shows the CLIP score of MSD and single-student methods. MSD4-ADM achieves a competitive CLIP score and beats the single student counterpart, DMD2. We believe the CLIP score can increase further if one trains on the LAION dataset [61] instead of the COYO dataset [8].
A.2 Consistency Distillation, Toy Experiments
To show the wider applicability of MSD, we apply another distillation method, Consistency Distillation [65], on the toy experiment setting in Sec. 5.1. Fig. 6 displays generated samples and the $\ell_{1}$ distance from teacher generation. While noting that consistency distillation achieves a weaker distillation of the teacher in general, we again observe better performance for more students. This indicates the generality of MSD.
[…]
Figure 6: A 2D toy model using consistency distillation. From left to right: teacher (multi-step) generation and student, one-step generation with $1$ and $8$ distilled students, the $\ell_{1}$ distance of generated samples between teacher and students. Takeaway: More students improve distillation quality on this easy-to-visualize setup with consistency distillation.
Appendix B Additional ablation studies
B.1 Training curves for Sec. 5.4
@@ Appendix B Additional ablation studies
B.2 The effect of paired dataset size on DMD
@@
B.3 Single-step distillation methods comparison
@@
Method on ImageNet Terminal FID Iterated Images
@@
Table 6 justifies our choice of DMD/DMD2 as our first-stage training without adversarial loss. Competitor methods either need larger training data size (EMD, SiD) or have worse quality (CTM and other CM-based methods). DMD/DMD2, on the other hand, strike a good balance. DMD exhibits more stability for MSD on ImageNet, whereas DMD2 performs better for SD v1.5, which leads to our respective choices. As mentioned in App. D, we used a smaller-sized paired dataset (10 0001000010\,00010 000 images) than the original DMD paper (25 0002500025\,00025 000 images) for ImageNet, which significantly accelerated convergence without impairing the final performance. Moreover, as in Sec. 4.2, the same paired dataset can be used for all students, eliminating potential additional computation.
B.4 More results on distilling into smaller students
@@
Figure 9: The effect of student size, number of classes covered, and training stage on the performance of a single smaller student. XS, XXS, and B refer to model sizes, with precise definitions provided in Tab. 7. Also, see special evaluation details in App. E.
In Sec. 5.2, we trained MSD4-ADM on smaller students to demonstrate the tradeoff between generation quality and speed. Here, we make a more comprehensive ablation study on the interplay between student size, number of classes covered, and training stage, with results in Fig. 9. We observe that generation quality increases with student size and decreases with more classes covered. MSD offers great flexibility for users to make these choices based on computational resources, generation quality, and inference speed requirements.
Appendix C Deployment suggestions
Here, we discuss some MSD deployment options for practitioners.
A naive option for deployment is to use increased GPU memory to host all models simultaneously. However, this is impractical and wasteful, as only a single student model needs to be used for each user request. In settings with less GPU memory than all students’ sum memory requirement, we must swap student models on and off GPUs. This incurs extra latency. However, in the few-GPU many-users setting, there are already prominent latency issues, such as users needing to queue for usage. In few-user settings, resources are likely being taken offline to save cost, and thus, there is start-up latency for fresh requests, too. Therefore, we argue that the more interesting setting is in large distributed deployment.
For settings with more GPU memory than all students’ sum memory requirements, we can distribute the student models among a cluster of GPUs (as one would the teacher) and route each generation request to the appropriate student node. The routing layer is lightweight compared to the inference cost, so we pay little for it.
If the data has been partitioned uniformly according to user demand, the incoming requests will be distributed uniformly among the student nodes. Therefore, we achieve equal throughput compared to the teacher without more overall model storage. However, finding such a partition is challenging, and user demand may change over time. This leaves finding the optimal allocation of resources to the student nodes an open problem. In practice, we expect that a reduced student model size would lead to an overall reduction in storage requirements compared to the teacher alone.
Appendix D Implementation Details
D.1 Toy Experiments
The real dataset is a mixture of Gaussians. The radius for the outer circle is $0.5$, the radius for the $8$ inner circles is $0.1$, and the standard deviation for each Gaussian (smallest circle) is $0.005$. The teacher is trained with EDM noise schedule, where use $(\sigma_{\min},\sigma_{\max})=(0.002,80)$ and discretized the noise schedule into $1000$ steps. We train the teacher for 100 000100000100\,000100 000 iterations with AdamW [38] optimizer, setting the learning rate at 1e-4, weight decay to $0.01$, and beta parameters to $(0.9,0.999)$. For distillation, we first generated a dataset of $1000$ pairs, then used DMD [76] to train $1$, $2$, $4$, and $8$ students, respectively, all for 200 000200000200\,000200 000 iterations, with reduced learning rate at 1e-7. We only sample the first $750$ of the $1000$ steps for distillation.
@@ D.1 Toy Experiments
D.2 ImageNet
D.2.1 Same-sized students
@@
For the ADM stage, we attach a prediction head to the middle block of the fakescore model. The prediction head consists of a stack of $4\times 4$ convolutions with a stride of $2$, group normalization, and SiLU activations. All feature maps are downsampled to $4\times 4$ resolution, followed by a single convolutional layer with a kernel size and stride of $4$. The final output linear layer maps the given vector to a scalar predicted probability. We load the best generator checkpoint from DM stage but re-initialize the fakescore and GAN classifier model from teacher weights, as we observed this leads to slightly better performance. We set the GAN generator loss weight to 3e-3 and the GAN discriminator loss weight to 1e-2 and reduce the learning rate to 5e-7. Each student is trained on $4$ A100 GPUs, with a total batch size of $192$, for 150 000150000150\,000150 000 iterations. This yields the MSD4-ADM checkpoint in Tab. 2.
D.2.2 Smaller students
@@
For the MSD4-ADM-S checkpoint in Tab. 2, we train the TSM stage using the model architecture S with the continuous EDM noise schedule with $(P_{\text{mean}},P_{\text{std}})=(-1.2,1.2)$ and the weighting $\lambda_{t}=(\sigma_{t}^{2}+0.5^{2})/(\sigma_{t}\cdot 0.5)^{2}$. We use a learning rate of 1e-4. Each student is trained on $4$ A100 GPUs, with a total batch size of $576$, for 400 000400000400\,000400 000 iterations. Then DM and ADM stages were trained using a total batch size of $160$, following otherwise the same setup as Sec. D.2.1.
For the ablation study in B.4, we instead train a common TSM stage for all students for computational efficiency. We train this common stage using $16$ A100 GPUs, with a total batch size of $3584$ and $4864$ for architecture XS and XXS, respectively. The DM and ADM stages are followed by specialized students with filtered data and $4$ A100 GPUs each, with a total batch size of $224$ and $256$ for architecture XS and XXS, respectively, and using the same learning rate of 2e-6.
D.3 SD v1.5
D.3.1 Same-sized students, CFG=1.75
Our SD v1.5 experiments setup closely follows DMD2 [75] paper. We distill our one-step generators from the SD v1.5 [54] model, using a classifier-free guidance scale of $1.75$ for the teacher model to obtain the best FID score. We use the first $5$M prompts from the COYO dataset [8] and the corresponding $5$M images for the GAN discriminator. We apply the DDIM noise schedule with $1000$ steps for sampling $t$. The weight $w_{t}$ in Eq. 2 is set to $\frac{\sigma_{t}^{2}}{\alpha_{t}}\frac{CS}{\|\bm{\mu}_{\text{teacher}}(\bm{x}_{t},t)-\bm{x}\|_{1}}$ where $S$ is the number of spatial locations and $C$ is the number of channels, and the weight $\lambda_{t}$ in Eq. 4 is set to $\alpha_{t}^{2}/\sigma_{t}^{2}$.
@@
For the ADM stage, we attach a prediction head that has the same architecture (though with a different input size) as Sec. D.2. We load both the best generator checkpoint and the corresponding fakescore checkpoint from DM stage. We set the GAN generator loss weight to 1e-3 and the GAN discriminator loss weight to 1e-2 and reduced the learning rate to 5e-7. Each student is trained on $32$ A100 GPUs, with a total batch size of $1024$, for $5000$ iterations. This yields the MSD4-ADM checkpoint in Tab. 2.
D.3.2 Same-sized students, CFG=8
@@ D.3.3 Smaller student, CFG=8
To create a subset of dog-related data, we first selected 1 210 000similar-toabsent1210000$\sim$\$1\,210\,000\$∼ 1 210 000 prompts in the COYO [8] dataset whose embeddings are closest to “a dog.” We then created an equal number of noise-image pairs from the SD v1.5 teacher using these prompts with CFG $=8$. We train the TSM stage using the model architecture S with the $1000$-step DDIM noise schedule and the weighting $\lambda_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$. We use a learning rate of 1e-4. We then continue to the DM stage with the paired regression loss, using a learning rate of 1e-5, and finally continue to the ADM stage using generated paired images as “real” images with a learning rate of 5e-7. We use $16$ A100 GPUs. The TSM stage is trained with a total batch size of $1536$ for 240 000240000240\,000240 000 iterations. The DM stage is trained with a total batch size of $512$ for both the paired and the fake images for 20 0002000020\,00020 000 iterations, and the ADM stage is trained with the same batch size for $6000$ iterations. This yields the checkpoint used to generate Fig. 2(c). Longer training and better tuning are again likely to improve the generation quality further.
Appendix E Evaluation Details
For zero-shot COCO evaluation, we use the exact setup as GigaGAN [77] and DMD2 [75]. Specifically, we generate 30 0003000030\,00030 000 images using the prompts provided by DMD2 code. We downsample the generated images using PIL to $256\times 256$ Lanczos resizer. We then use the clean-fid [49] to compute the FID score between generated images and 40 5044050440\,50440 504 real images from the COCO 2014 validation dataset. Additionally, we use the OpenCLIP-G backbone to compute the CLIP score. For ImageNet, we generate 50 0005000050\,00050 000 images and calculate FID using EDM [25] evaluation code. When selecting the best checkpoints for partitioned students, the same procedure is applied only for prompts/classes within respective partitions. For the ablation study in Sec. B.4, 10 0001000010\,00010 000 images are generated for only the first $10$ classes for an apple-to-apple comparison.
Appendix F Additional qualitative results
F.1 Additional Imagenet-64$\times$64 results